At What temperature will a sucrose solution boy if it contains 1.55 Mol of sucrose in 600 mL of water? Kb of water is 0.51 °C/m


1
Expert's answer
2021-10-14T04:03:45-0400

"ΔT_b = K_b \times m \\

= 0.51 \times \frac{1.55 \times 1000}{600} = 1.3175 \\

ΔT_b = T_1-T_0 \\

T_1 = ΔT_b + T_0 \\

= 1.32 + 100 = 101.32 \;°C"

Answer: 101.32 °C